OpenStudy (anonymous):

Sodium metal used in photoelectric experiment, illuminated with light wavelength 420nm. Stopping potential is .65volts. a. What is sodium's work function? b. What wavelength would produce a stopping potential of 1.69V with the same setup?

OpenStudy (anonymous):

The KE of the electron=exV (V=stopping potential, e=charge of electron=1.6/10^19 C) From Einstein photoelectric equation hc/l=work function + KE(l=wavelength of incident light,c=speed of light,h=Planck's constant=6.62/10^32 Js) b. Say the required wavength is l' Again hc/l'=work function(known from above)+ex1.69